# Density control of multi-agent swarms via bio-inspired leader-follower plasticity

### v1

@@ Abstract
The design of control systems for the spatial self-organization of mobile agents is an open challenge across several engineering domains, including swarm robotics and synthetic biology. Here, we propose a bio-inspired leader-follower solution, which is aware of energy constraints of mobile agents and is apt to deal with large swarms. Akin to many natural systems, control objectives are formulated for the entire collective, and leaders and followers are allowed to plastically switch their role in time. We frame a density control problem, modeling the agents’ population via a system of nonlinear partial differential equations. This approach allows for a compact description that inherently avoids the curse of dimensionality and improves analytical tractability. We derive analytical guarantees for the existence of desired steady-state solutions and their local stability for one-dimensional and higher-dimensional problems. We numerically validate our control methodology, offering support to the effectiveness, robustness, and versatility of our proposed bio-inspired control strategy.
@@
thanks: Corresponding authors Mario di Bernardo and Maurizio Porfiri.
@@ 1 Introduction
Across a wide range of applications, from swarm robotics [1] to collective construction [2], environmental management [3], synthetic biology [4], and search-and-rescue operations [5], engineers are to design the spatial displacement of large swarms. Leader-follower strategies [6] are commonly used toward this objective, usually assuming that a population of controller agents (leaders) is given the task of inducing a desired behavior in another population (followers). The implementation of these leader-follower strategies is often informed by our understanding of the mechanisms underpinning collective behavior of various animal species, such as fish [7, 8], birds [9], and humans [10, 11]. Despite considerable success, existing engineered systems fail to capture the richness of their biological counterparts in two key aspects.
First, leaders and followers are part of a fixed hierarchy, which does not capture role switching due to behavioral plasticity. Behavioral plasticity is the reversible biological mechanism behind changes in the behavior of organisms due to internal or environmental stimuli [12]. For instance, in fish schools behavioral plasticity arises when followers become leaders as they access novel information about predators and food locations [13, 14]. Likewise, in flocks of migrating birds, it occurs as leaders step down from frontal positions in the formation to recover energy [15]. Similar observations are gathered in human groups performing collective tasks, such as games [16] and team sports [17].
Second, leaders in engineered systems are typically viewed as fixed reference signals in open-loop, not receptive to the unfolding collective dynamics [18]. Even when a feedback loop is included in the design of the leader-follower control strategy, its goal is only to steer the behavior of followers without taking into account the entire collective [19]. In biological systems, the behavior of leaders adapts over time in response to collective needs; for example, foraging or predator avoidance are key to survival and reproduction of both leaders and followers [20].
Addressing both of these gaps would benefit several application areas related to spatial organization of large swarms. For example, a recent perspective [21] has noted that the implementation of plasticity in robotic swarms may be a viable option to face the high unpredictability of the real world and grant agents the necessary flexibility needed to perform a wide spectrum of tasks. Enabling spontaneous roles’ switching between leaders and followers may be conducive to an overall reduction of control efforts during task execution, thereby containing energy costs for the swarm. Such a benefit may further extend to the domain of synthetic biology, where there is a need for control schemes that could regulate the composition of microbial consortia composed to guarantee efficient labor division [22]. Here, we develop a continuum model for a large population of behaviorally plastic leaders and followers agents solving a density control problem.
@@
Mathematically, the control of large swarms introduces several challenges: ($i$) the state space grows exponentially with the number of agents; ($ii$) communication graphs may be time-varying and constrained; and ($iii$) individual agent dynamics are coupled through nonlocal interaction terms. Continuum approaches using partial differential equations (PDEs) offer a promising avenue to circumvent these difficulties by modeling agent densities rather than individual states, thus achieving dimension reduction while preserving essential collective dynamics [23, 24, 25]. Mean-field models of interacting populations of mobile agents have gained substantial momentum over the last decade [26, 27, 28], as they offer compact formulations to increase computational and analytical tractability. Bio-inspired switching mechanisms between leaders and followers have been explored in [29, 30, 31, 32, 33] for modeling purposes, but none of these efforts has presented feedback control actions to induce desired collective behaviors. A mean-field optimal control problem with transient leadership has been formulated in [34], but, due to the absence of closed form feedback solutions, it can only be numerically approximated, yielding limited insight that may inform control design in realistic settings. Related work on leader selection [35, 36] and switching mechanisms [37] exists, but these approaches either maintain fixed population assignments or focus on synchronization rather than spatial density control.
@@
We draw insight from the literature on reacting mixtures [38], used to describe blood flows [39] and tumors growth [40]. In particular, we model plasticity as a chemical reaction taking place between two fluids, associated with the continuum description of leaders and followers. We assume that plasticity is not common to all agents, where some of them may not be allowed switching their role. Whether or not an agent is a leader or a follower is not distinguishable by the rest of the group. We derive conditions for the existence and stability of the solution, in terms of key model parameters that can be part of engineering design, such as the fraction of non-plastic agents, interaction kernels, and desired densities. Our approach enables us to provide explicit feedback control laws that guarantee exponential convergence, stability analysis for the coupled PDE system, and feasibility conditions through adaptive leader selection.
The main control-theoretic contributions of the paper are as follows: ($i$) We derive explicit feedback control laws for density regulation in heterogeneous swarms with adaptive leader-follower role assignment, providing exponential convergence guarantees. We include plastic leaders and followers, which can dynamically exchange roles, and non-plastic followers, which cannot switch to leadership roles; ($ii$) we establish necessary and sufficient conditions for the existence and local stability of desired equilibrium density distributions under role-switching dynamics; ($iii$) we extensively validate our strategy through numerical experiments, demonstrating that adaptive role switching enhances system robustness to parametric uncertainties compared to fixed-hierarchy approaches.
@@
The rest of the paper is organized as follows. In Sec. 2, we present the problem statement for a one-dimensional (1D) scenario. The proposed control strategy is formulated in Sec. 3 and its numerical validation is detailed in Sec. 4. An agent-based model that supports the use of our continuum approach is presented in Sec. 5. The extension to higher dimensions and the corresponding numerical validation are expounded in Sec. 6. Section 7 concludes the manuscript summarizing our main findings and proposing avenues of future research.
@@ 2 Problem Statement
Our control objective is to steer the spatial distribution of the entire collective toward a desired configuration through coordinated leader actions, while allowing adaptive role assignment within the swarm. We adopt a continuum approach to describe the densities of the agents. The mathematical formulation consists of three coupled convection-diffusion equations on the unit circle $\mathcal{S}=[-\pi,\pi]$,
$\rho^{L}_{t}(x,t)+\left[\rho^{L}(x,t)u(x,t)\right]_{x}+$ $\left[\rho^{L}(x,t)(f*\rho)(x,t)\right]_{x}$
$=D$ $\rho^{L}_{xx}(x,t)+q(x,t),$ (1a)
$\rho^{F}_{t}(x,t)+\big{[}\rho^{F}(x,t)(f*\rho)(x,t)$ $\big{]}_{x}=D\rho^{F}_{xx}(x,t)$
$\qquad\quad-q(x,t),$ (1b)
$\eta^{F}_{t}(x,t)+\big{[}\eta^{F}(x,t)(f*\rho)(x,t)$ $\big{]}_{x}=D\eta^{F}_{xx}(x,t).$ (1c)
Here, $(\cdot)_{t}$ and $(\cdot)_{x}$ indicate partial derivatives with respect to time and spatial coordinates, respectively. These equations model the spatio-temporal dynamics of the density of three subsets of agents, respectively
  • leaders (whose density is $\rho^{L}:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$) that ($i$) react to the rest of the group, ($ii$) are controlled through the periodic velocity field $u:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$, and ($iii$) may become followers due to the reacting mechanism $q:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$;
  • plastic followers (whose density is $\rho^{F}:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$) that ($i$) react to the rest of the population, and ($ii$) may become leaders due to the reacting mechanism $q$;
  • non-plastic followers (whose density is $\eta^{F}:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$) that only react to the rest of the population but cannot switch their role to leaders.
We consider all-to-all inter- and intra-population interactions taking place through a periodic interaction kernel $f$ (soft-core and vanishing with distance). We model these interactions through a cross-convectional non-local term involving $\rho=\rho^{L}+\rho^{F}+\eta^{F}$. To account for the presence of noise affecting the behavior of any agent in the real world, we include a diffusion process, weighted by the diffusion coefficient $D$.
By summing (1a), (1b) and (1c), we obtain the dynamics of the density of the whole population
$\rho_{t}(x,t)+\left[\rho^{L}(x,t)u(x,t)\right]_{x}+\left[\rho(x,t)(f*\rho)(x,t)\right]_{x}=\\ =D\rho_{xx}(x,t),$ (2)
which is independent of the reacting term $q$. By imposing periodic boundary conditions for (1a), (1b) and (1c),
@@
we ensure that the total population mass is conserved. In particular, we have
$\left(\int_{\mathcal{S}}\rho(x,t)\,\mathrm{d}x\right)_{t}=-\left[\rho^{L}(x,t)u(x,t)\right]_{-\pi}^{\pi}\\ -\left[\rho(x,t)(f*\rho)(x,t)\right]_{-\pi}^{\pi}+D\left[\rho_{x}(x,t)\right]_{-\pi}^{\pi}=0,$ (4)
because of periodicity. Notice that (1c) satisfies mass conservation, but (1a) and (1b) do not, due to the reacting term $q$. If $\int_{\mathcal{S}}q\,\mathrm{d}x$ equals 0, there is no net mass transfer between leaders and followers.
Without loss of generality, we normalize the total mass to one by setting
@@
where $M^{L}$ is the leaders’ mass, $M^{F}$ is the followers’ mass, and $\Phi^{F}$ is the constant mass of non-plastic followers. We define the fraction of the population that is allowed to switch role, $p$, as
@@ 3.1 Design of the leaders’ velocity field
We define the error function for the density of the entire collective,
@@
$e_{t}(x,t)=\left[\rho^{L}(x,t)u(x,t)\right]_{x}+\left[\rho(x,t)(f*\rho)(x,t)\right]_{x}\\ -D\rho_{xx}(x,t),$ (11)
@@
Assume $\rho^{L}>0$ for any $x\in\mathcal{S}$ and $t\in\mathbb{R}_{\geq 0}$. Choosing the control input $u$ in (1a) according to
$\left[\rho^{L}(x,t)u(x,t)\right]_{x}=-Ke(x,t)-\left[\rho(x,t)(f*\rho)(x,t)\right]_{x}\\ +D\rho_{xx}(x,t),$ (12)
where $K>0$ is a control gain, the error converges globally and exponentially to 0 point-wise in $\mathcal{S}$, that is,
$e(x,t)=e(x,0)\,\mathrm{exp}(-Kt).$ (13)
@@
The control input $u$ can be found by spatial integration of (12), as
$u(x,t)=\frac{1}{\rho^{L}(x,t)}\bigg{[}-K\int e(x,t)\,\mathrm{d}x$
$-\rho(x,t)(f*\rho)(x,t)$ $+D\rho_{x}(x)\bigg{]},$ (15)
which is well-defined only for $\rho^{L}>0$. Note that $u$ can be shown to be periodic (see Corollary 1 in [19]).
@@
Under the hypothesis of Theorem 1 and using (13), we establish the following closed-form expression for the density of the collective:
@@ 3.2 Design of the reacting term
We achieve these goals by choosing the reacting term $q$ as
$q(x,t)=\frac{1}{2}\left[\rho^{L}(x,t)u(x,t)\right]_{x}+\frac{1}{2}\left[\rho^{*}(x,t)(f*\rho)(x,t)\right]_{x}\\ -\frac{D}{2}\rho^{*}_{xx}(x,t)+g(x,t),$ (17)
@@
and $g$ obeys to the mass action law
@@
$\rho^{*}_{t}(x,t)=2q(x,t)-\left[\rho^{L}(x,t)u(x,t)\right]_{x}\\ -\left[\rho^{*}(x,t)(f*\rho)(x,t)\right]_{x}+D\rho^{*}_{xx}(x,t).$ (22)
@@
$\rho^{L}(x,t)$ $=\frac{1}{2}\left[\rho(x,t)+\rho^{*}(x,t)-\eta^{F}(x,t)\right],$ (23a)
$\rho^{F}(x,t)$ $=\frac{1}{2}\left[\rho(x,t)-\rho^{*}(x,t)-\eta^{F}(x,t)\right].$ (23b)
@@
$\rho^{*}_{t}(x,t)=-a\,\rho^{*}(x,t)+b\,\rho(x,t)-b\,\eta^{F}(x,t),$ (24)
where
$a$ $:=K_{FL}+K_{LF},$ (25a)
$b$ $:=K_{FL}-K_{LF}.$ (25b)
Under the control action discussed in Theorem 1, $\bar{\rho}$ is a steady-state solution for (2), so that we look for steady-state solutions of (24) and (1c). We start by considering (1c) with $\eta^{F}_{t}=0$, $\eta^{F}(x,t)=\bar{\eta}^{F}(x)$, and $\rho(x,t)=\bar{\rho}(x)$, which gives
$D\bar{\eta}^{F}_{xx}(x)-\left[\bar{\eta}^{F}(x)(f*\bar{\rho})(x)\right]_{x}=0.$ (26)
@@
$\bar{\rho}^{L}(x)=\frac{1}{2}\left[\bar{\rho}(x)\left(1+\frac{b}{a}\right)-\bar{\eta}^{F}(x)\left(1+\frac{b}{a}\right)\right],$ (29a)
$\bar{\rho}^{F}(x)=\frac{1}{2}\left[\bar{\rho}(x)\left(1-\frac{b}{a}\right)-\bar{\eta}^{F}(x)\left(1-\frac{b}{a}\right)\right].$ (29b)
@@
($\implies$) The existence of a steady-state solution for (1) with $\rho^{L},\rho^{F}>0$ and $\bar{\eta}^{F}$ implies that
$\bar{\rho}^{L}(x)+\bar{\rho}^{F}(x)>0,\;\forall\,x\in\mathcal{S}.$ (31)
By adding and subtracting $\bar{\eta}^{F}$, we obtain
$\bar{\rho}^{L}(x)+\bar{\rho}^{F}(x)+\bar{\eta}^{F}(x)>\bar{\eta}^{F}(x),\;\forall\,x\in\mathcal{S},$ (32)
@@
Theorem 4 gives conditions about the minimum fraction of agents that can switch role such that $\bar{\rho}$ can be a meaningful steady-state solution for (1), that is (5) and (8) hold with $\rho^{L}>0$, $\rho^{F},\eta^{F}\geq 0$. Hence, it consists in an existence result for the solutions of the problem given in Sec. 2.
@@
The requirement in (9) can be ensured by choosing $K_{LF}$ and $K_{FL}$ in (19).
@@
$\frac{\int_{\mathcal{S}}\bar{\rho}^{L}(x)\,\mathrm{d}x}{\int_{\mathcal{S}}\bar{\rho}^{F}(x)\,\mathrm{d}x}=\frac{K_{FL}}{K_{LF}}.$ (34)
Hence, by appropriately choosing the reacting rates $K_{LF}$ and $K_{FL}$, we fulfill (9).∎
@@
where we fix mean $\mu=0$ and concentration coefficient $k=1$. $Z$ is chosen so that $\bar{\rho}$ integrates to 1. Using (20), we establish that the fraction of agents that are allowed to switch the role should be larger than $\hat{p}\approx 0.15$ for the desired collective density to be feasible.

### v2

@@ Abstract
The design of control systems for the spatial self-organization of mobile agents is an open challenge across several engineering domains, including swarm robotics and synthetic biology. Here, we propose a bio-inspired leader-follower solution, which is aware of energy constraints of mobile agents and is apt to deal with large swarms. Akin to many natural systems, control objectives are formulated for the entire collective, and leaders and followers are allowed to plastically switch their role in time. We frame a density control problem, modeling the agents’ population via a system of nonlinear partial differential equations. This approach allows for a compact description that inherently avoids the curse of dimensionality and improves analytical tractability. We derive analytical guarantees for the existence of desired steady-state solutions and their global stability for one-dimensional and higher-dimensional problems. We numerically validate our control methodology, offering support to the effectiveness, robustness, and versatility of our proposed bio-inspired control strategy.
@@
thanks: This paper was not presented at any IFAC meeting. Corresponding authors Mario di Bernardo and Maurizio Porfiri.
@@ 1 Introduction
Across applications, from swarm robotics [1] to search-and-rescue operations [2], leader-follower strategies [3] are used to design the spatial displacement of large swarms, in which controller agents (leaders) are to induce a desired spatial behavior of another group of agents (followers). While often informed by biological systems like fish [4, 5] and birds [6], engineered systems fail to capture two key aspects of their biological counterparts.
First, leaders and followers maintain fixed hierarchies [7] missing behavioral plasticity, that is, the reversible mechanism behind role changes in biological systems due to internal or environmental stimuli [8]. In fish schools, plasticity arises when followers access novel information about predators or food [9]; likewise, in migrating birds, leaders step down from frontal positions in the flock to recover energy [10]. Second, engineered leaders typically act as fixed open-loop signals, unreceptive to the unfolding collective dynamics [11], whereas biological leaders adapt based on collective needs critical to survival, such as in foraging and predator avoidance [12]. Capturing both these aspects would benefit applications in spatial organization of large swarms, granting flexibility to address multiple tasks at lower control efforts and energy costs [13].
@@
Mathematically, the control of large swarms poses several challenges: ($i$) the state space grows exponentially with the number of agents; ($ii$) communication graphs may be time-varying and constrained; and (iiiiii) individual agent dynamics are non-locally coupled. Continuum approaches using partial differential equations (PDEs) offer a promising avenue to circumvent these difficulties by modeling agent densities rather than individual states, achieving dimension reduction preserving essential collective dynamics [14, 15, 16]. Mean-field models of interacting populations of agents have gained substantial momentum in the last decade [17, 18, 19], as they offer compact formulations to increase computational and analytical tractability. Bio-inspired switching mechanisms between leaders and followers have been explored in [20, 21, 22, 23, 24] for modeling purposes, lacking the presentation of feedback control actions to induce desired collective behaviors. A mean-field optimal control problem with transient leadership has been formulated in [25], but, due to the absence of closed form feedback solutions, it can only be numerically approximated, yielding limited insight informing control design in realistic settings. Related work on leader selection [26, 27] and switching mechanisms [28] exists, but these approaches either maintain fixed population assignments or focus on synchronization rather than spatial density control.
@@
Here, we develop a continuum model for a large population of behaviorally plastic leaders and followers solving a density control problem. We draw insight from the literature on reacting mixtures [29], used to describe blood flows [30] and tumors growth [31]. In particular, we model plasticity as a chemical reaction taking place between two fluids, associated with the continuum description of leaders and followers. We assume that plasticity is not common to all agents, where some of them may not be allowed switching their role. Whether or not an agent is a leader or a follower is not distinguishable by the rest of the group. We derive conditions for the existence and stability of the solution, in terms of key parameters that can be part of engineering design. Our approach provides explicit feedback control laws that guarantee convergence of the coupled PDE system through adaptive leader selection.
@@
The rest of the paper is organized as follows. In Sec. 2, we present the problem statement for a one-dimensional (1D) scenario. The control strategy is formulated in Sec. 3 and validated in Sec. 4. An agent-based model supporting our continuum approach is presented in Sec. 5. The extension to higher dimensions and its validation are expounded in Sec. 6. Section 7 concludes the manuscript.
@@ 2 Problem Statement
Our control objective is to steer the spatial distribution of a large-scale swarm towards a desired configuration through coordinated leader actions and adaptive role assignment. We adopt a continuum approach consisting of coupled convection-diffusion-reaction equations on the unit circle $\mathcal{S}=[-\pi,\pi]$ [^1] For the sake of brevity, throughout the paper, explicit dependencies on spatial coordinate $x$ and time $t$ are omitted when unambiguous.
$\rho^{L}_{t}+\left(\rho^{L}u\right)_{x}+\left[\rho^{L}(f*\rho)\right]_{x}$ $=D\rho^{L}_{xx}+q,$ (1a)
$\rho^{F}_{t}+\left[\rho^{F}(f*\rho)\right]_{x}$ $=D\rho^{F}_{xx}-q,$ (1b)
$\eta^{F}_{t}+\left[\eta^{F}(f*\rho)\right]_{x}$ $=D\eta^{F}_{xx},$ (1c)
where $(\cdot)_{t}$ and $(\cdot)_{x}$ denote partial derivatives. The densities $\rho^{L}$, $\rho^{F}$, $\eta^{F}:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}_{\geq 0}$ are associated with three agents’ subsets: leaders controlled through the periodic velocity field $u:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$ that may become followers; plastic followers that may become leaders; and non-plastic followers that cannot be engaged in roles’ switching. Plasticity between leaders and followers take place through the reaction term $q:\mathcal{S}\times\mathbb{R}_{\geq 0}\to\mathbb{R}$; inter- and intra-population interactions are modeled through a cross-convectional non-local term involving the interaction kernel $f$ (soft-core and odd) and the swarm density $\rho=\rho^{L}+\rho^{F}+\eta^{F}$; agent-level stochasticity is captured by diffusion weighted by the positive coefficient $D$.
Summing eqs. (1) yields the total population density dynamics
$\rho_{t}+\left(\rho^{L}u\right)_{x}+\left[\rho(f*\rho)\right]_{x}=D\rho_{xx},$ (2)
which is independent of $q$. Periodic boundary conditions,
@@
ensure mass conservation, that is $\left(\int_{\mathcal{S}}\rho\,\mathrm{d}x\right)_{t}=0$ (exploiting periodicity of the functions). Note that (1c) satisfies mass conservation on its own, but (1a) and (1b) do not, due to the reaction $q$. If $\int_{\mathcal{S}}q\,\mathrm{d}x\,{=0}$, there is no net mass transfer between leaders and followers. Eqs.  (1) are also complemented by the initial conditions
$\rho^{i}(x,0)=\rho^{i}_{0}(x),\,i=L,F,\quad\eta^{F}(x,0)=\eta^{F}_{0}(x).$ (4)
Preserving generality, we normalize the total mass to one
@@
where $M^{L}$ and $M^{F}$ are the leaders’ and followers’ mass, and $\Phi^{F}$ is the constant mass of non-plastic followers. The fraction of the population allowed to switch role, $p$, is
@@ 3.1 Design of the leaders’ velocity field
We define the error function for the the collective,
@@
$e_{t}=\left(\rho^{L}u\right)_{x}+\left[\rho(f*\rho)\right]_{x}-D\rho_{xx},$ (10)
@@
Assume $\rho_{L}>0$ for any $x\in\mathcal{S}$ and $t\in\mathbb{R}_{\geq 0}$. Choosing the control input as
$u=\frac{1}{\rho^{L}}\left[-K\int e\,dx-\rho(f*\rho)+D\rho_{x}\right],$ (11)
where $K>0$, the error converges to 0 globally and exponentially, that is,
$e(x,t)=e(x,0)\exp(-Kt).$ (12)
@@
The control input $u$ is well-defined only for $\rho^{L}>0$. Note that $u$ can be shown to be periodic (see Corollary 1 in [7]).
@@
From (12), we establish the following closed-form expression for the density of the collective
@@ 3.2 Design of the reacting term
We achieve these goals choosing the reacting term $q$ as
$q=\frac{1}{2}\left(\rho^{L}u\right)_{x}+\frac{1}{2}\left[\rho^{*}(f*\rho)\right]_{x}-\frac{D}{2}\rho^{*}_{xx}+g,$ (15)
@@
and $g$ obeys the mass action law
@@
Remark 4.
The reaction term $q$ can be factorized as
This represents a mass action law whose non-constant reaction rates are given by (using (15) and (11))
$\kappa_{FL}$ $=K_{FL}-\frac{Ke}{2\rho^{F}}+\frac{D\rho_{xx}}{2\rho^{F}}-\frac{1}{2\rho^{F}}\left[\rho(f*\rho)\right]_{x},$ (19a)
$\kappa_{LF}$ $=K_{LF}+\frac{D\rho^{*}_{xx}}{2\rho^{L}}-\frac{1}{2\rho^{L}}\left[\rho^{*}(f*\rho)\right]_{x}.$ (19b)
Note that this factorization is not unique.
@@
$\rho^{*}_{t}=2q-\left(\rho^{L}u\right)_{x}-\left[\rho^{*}(f*\rho)\right]_{x}+D\rho^{*}_{xx}.$ (22)
@@
$\rho^{L}$ $=(1/2)\left[\rho+\rho^{*}-\eta^{F}\right],$ (23a)
$\rho^{F}$ $=(1/2)\left[\rho-\rho^{*}-\eta^{F}\right].$ (23b)
@@
$\rho^{*}_{t}=-a\,\rho^{*}+b\,\rho-b\,\eta^{F},$ (24)
where $a$ and $b$ are the constants defined as
$a:=K_{FL}+K_{LF},\quad b:=K_{FL}-K_{LF}.$ (25)
Compared to the complex dynamics of $\rho^{L}$ and $\rho^{F}$, the dynamics of $\rho$ and $\rho^{*}$ is linear. Moreover, under the control in Theorem 1, $\bar{\rho}$ is a steady-state solution for (2), so that we look for steady-state solutions of (24) and (1c). We start by considering (1c) with $\eta^{F}_{t}=0$, $\eta^{F}(x,t)=\bar{\eta}^{F}(x)$, and $\rho(x,t)=\bar{\rho}(x)$, which gives
$D\bar{\eta}^{F}_{xx}-\left[\bar{\eta}^{F}(f*\bar{\rho})\right]_{x}=0.$ (26)
@@
$\bar{\rho}^{L}$ $=\frac{1}{2}\left[\bar{\rho}\left(1+b/a\right)-\bar{\eta}^{F}\left(1+b/a\right)\right],$ (29a)
$\bar{\rho}^{F}$ $=\frac{1}{2}\left[\bar{\rho}\left(1-b/a\right)-\bar{\eta}^{F}\left(1-b/a\right)\right].$ (29b)
@@
($\implies$) The existence of a steady-state solution for (1) with $\rho^{L},\rho^{F}>0$ and $\bar{\eta}^{F}\geq 0$ implies that $\bar{\rho}^{L}+\bar{\rho}^{F}>0$ $\forall x\in\mathcal{S}$. By adding and subtracting $\bar{\eta}^{F}$, we obtain
$\bar{\rho}^{L}+\bar{\rho}^{F}+\bar{\eta}^{F}>\bar{\eta}^{F},\;\forall\,x\in\mathcal{S},$ (31)
@@
Theorem 5 gives conditions about the minimum fraction of agents that can switch role such that $\bar{\rho}$ can be a meaningful steady-state solution for (1), that is, (5) and (7) hold with $\rho^{L}>0$, $\rho^{F},\eta^{F}\geq 0$.
@@
The requirement in (8) can be ensured by appropriately choosing $K_{LF}$ and $K_{FL}$ in (17).
@@
$\frac{\int_{\mathcal{S}}\bar{\rho}^{L}(x)\,\mathrm{d}x}{\int_{\mathcal{S}}\bar{\rho}^{F}(x)\,\mathrm{d}x}{=\frac{1+b/a}{1-b/a}=\frac{a+b}{a-b}}=\frac{K_{FL}}{K_{LF}}.$ (33)
Hence, by appropriately choosing the reaction rates $K_{LF}$ and $K_{FL}$, we fulfill (8).∎
Remark 8.
The control action embedded in $u$ (11) and $q$ (15) assumes full knowledge of the swarm dynamics. This easily accommodates swarm robotics scenarios, in which centralized sensing is in place. Moreover, enabling agents’ communication to estimate densities allows for fully distributed implementations [32, 33].
@@
where we fix mean $\mu=0$ and concentration coefficient $k=1$. $Z$ is chosen so that $\bar{\rho}$ integrates to 1. Using (20), we establish that the fraction of agents allowed to switch role should be larger than $\hat{p}\approx 0.15$ for the desired collective density to be non-negative and integrating to 1.